ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 225, 2018

Aizenbud, A. and D. Gourevitch Invariant generalized functions supported on an orbit

411

Anquela, J. A., T. Cortés and I. Shestakov Commuting U-operators and nondegenerate imbeddings of Jordan systems

871

Bhowmik, M. and S. Sen Uncertainty principles of Ingham and Paley–Wiener on semisimple Lie groups

193

Boyle, M., T. M. Carlsen and S. Eilers Flow equivalence of sofic shifts

111

Brendle, J., A. Brooke-Taylor, S.-D. Friedman and D. C. Montoya Cichoń’s diagram for uncountable cardinals

959

Brooke-Taylor, A. See J. Brendle

Burness, T. C. and M. Giudici Locally elusive classical groups  343

Carlsen, T. M. See M. Boyle Cortés, T. See J. A. Anquela

Dubouloz, A. and T. Kishimoto Cylinders in del Pezzo fibrations  797

Duminil-Copin, H., M. R. Hilário, G. Kozma and V. Sidoravicius Brochette percolation

479

Eilers, S. See M. Boyle Feigon, B., K. Martin and D. Whitehouse Periods and nonvan-

ishing of central L-values for GL(2n)

223

Fisher, D., M. Larsen, R. Spatzier and M. Stover Character varieties and actions on products of trees

889

Friedman, S.-D. See J. Brendle Gaspar-Arreola, M. Á., F. Hernández-Hernández and M.

Hrušák Scattered spaces from weak diamonds

427

Gerlach, M. Convergence of dynamics and the Perron–Frobenius operator

451

Giudici, M. See T. C. Burness Godoy, Y. and M. Salvai Polar factorization of conformal and pro-

jective maps of the sphere in the sense of optimal mass transport

465

Gourevitch, D. See A. Aizenbud

Hamenstädt, U. Ergodicity of the absolute period foliation  661

Hernández-Hernández, F. See M. Á. Gaspar-Arreola Hilário, M. R. See H. Duminil-Copin Hrušák, M. See M. Á. Gaspar-Arreola Hubert, P., S. Lelièvre, L. Marchese and C. Ulcigrai The

Lagrange spectrum of some square-tiled surfaces

553

Iommi, G., F. Riquelme and A. Velozo Entropy in the cusp and phase transitions for geodesic flows

609

Izhakian, Z., M. Knebusch and L. Rowen Decompositions of modules lacking zero sums

503

Jiang, D. and C. Wu Periods and (χ, b)-factors of cuspidal automorphic forms of symplectic groups

267

Keller, C., S. Smorodinsky and G. Tardos Improved bounds on the Hadwiger–Debrunner numbers

925

Kindler, G., N. Kirshner and R. O’Donnell Gaussian noise sensitivity and Fourier tails

71

Kionke, S. and M. Vannacci Positively finitely related profinite groups

743

Kirshner, N. See G. Kindler Kishimoto, T. See A. Dubouloz Knebusch, M. See Z. Izhakian Kobayashi, Y. and N. Tanaka Well-posedness for mutational equa-

tions under a general type of dissipativity conditions

1

Kozma, G. See H. Duminil-Copin Larsen, M. See D. Fisher Lelièvre, S. See P. Hubert Levitt, J. and M. Yakimov Quantized Weyl algebras at roots of

unity

681

Lu, Y. Global weak solutions for the chromatography system  721

Marchese, L. See P. Hubert Martin, K. See B. Feigon Montoya, D. C. See J. Brendle Nica, B. On bounded elementary generation for SL_{n} over polynomial

rings

403

Nogueira, J. M. Knot exteriors with all possible meridional essential surfaces

909

O’Donnell, R. See G. Kindler Patel, S. P. and D. Prasad Restriction of representations of meta-

plectic GL_{2}(F) to tori

525

Popov, A. I. Existence of indecomposable members in indecomposable spaces of operators

947

Prasad, D. See S. P. Patel Raghavan, D. and S. Todorcevic Suslin trees, the bounding num-

ber, and partition relations

771

Riquelme, F. See G. Iommi Rowen, L. See Z. Izhakian Salvai, M. See Y. Godoy Schroeter, F. and E. Shustin Refined elliptic tropical enumerative

invariants

817

Sen, S. See M. Bhowmik Shestakov, I. See J. A. Anquela Shustin, E. See F. Schroeter Sidoravicius, V. See H. Duminil-Copin Smorodinsky, S. See C. Keller Spatzier, R. See D. Fisher Stover, M. See D. Fisher

Sundaram, S. On conjugacy classes of S_{n} containing all irreducibles  321

Tanaka, N. See Y. Kobayashi Tardos, G. See C. Keller Todorcevic, S. See D. Raghavan Treviño, R. Flat surfaces, Bratteli diagrams and unique ergodicity

à la Masur

35

Ulcigrai, C. See P. Hubert Vannacci, M. See S. Kionke Velozo, A. See G. Iommi

Wang, Z. J. Cocycle rigidity of abelian partially hyperbolic actions  147

Whitehouse, D. See B. Feigon Wu, C. See D. Jiang Yakimov, M. See J. Levitt